If a half of a number is decreased by 7 then doubled, the result is the same as when the number is increased by 20. What is the number?
Answer by ankor@dixie-net.com(22740)

If a half of a number is decreased by 7 then doubled, the result is the same as when the number is increased by 20. What is the number?
:
2(.5x-7) = x + 20
x - 14 = x + 20
you can see right here that this is not possible